# Visualized Text-to-Image Retrieval

### v1

@@ 4.1 Experimental Setup
We evaluate on three challenging benchmarks: (1) INQUIRE-Rerank (Vendrow et al., 2024), a T2I retrieval benchmark requiring accurate knowledge of species appearance and behavior. We perform additional filtering to remove overly simple queries and call the resulting dataset INQUIRE-Rerank-Hard. (2) Visual-RAG (Wu et al., 2025), a T2I retrieval and VQA benchmark featuring visual knowledge intensive questions on features of natural species that are not commonly documented in text corpus. (3) Visual-RAG-ME, a new benchmark we introduce featuring queries that compare the same visual feature across multiple entities. We present the benchmark details in Appendix A.
@@
For all the three benchmarks, we evaluate T2I retrieval with Recall@$k$ and NDCG@$k$ with $k$ = 1, 10, 30. For Visual-RAG and Visual-RAG-ME, we additionally use an LLM judge to evaluate the end-to-end VQA accuracy, following Wu et al. (2025).
For the experiments presented in the main text, we use CLIP and E5-V as the retriever, GPT-4o (OpenAI, 2023b) as the downstream reader, and gpt-image-1 (OpenAI, 2025) as the T2I Model to generate $m=3$ images. We analyze more model choices in Appendix B and present all the prompts and other hyperparameters in Appendix C.
4.2 Retrieval Performance
Table 1 summarizes retrieval performance across all benchmarks and retrievers. We compare four strategies: using the original textual query, applying query expansion via an LLM, our proposed VisRet, and VisRet with only a single generated image ("-multi-image"). Across all datasets, VisRet consistently outperforms both the original query and query expansion baselines by a large margin. When CLIP is used as the retriever, VisRet outperforms the original query and LLM-based rephrase by 32.7% and 15.6% relatively higher NDCG@10 over three benchmarks. Similar trends hold when E5-V is used as the retriever, exhibiting 24.5% and 12.4% performance gain in NDCG@10. Further, using only one generated image as the query only slightly harms the performance, indicating the flexibility of VisRet. Table 2 presents several examples to demonstrate how the visualization step successfully captures subtle visual semantics implied by the original text query. We present more analyses on the T2I model and the rephrase LLM in Section B.1 and Section B.2.
@@ 4.2 Retrieval Performance
4.3 Downstream QA Performance
@@
To assess the utility of VisRet in real-world applications, we evaluate its downstream VQA accuracy in a RAG pipeline. We compare three settings: (1) using only the model’s internal knowledge, (2) RAG with original text query-based retrieval and (3) RAG with VisRet. Figure 2 shows the QA accuracy on Visual-RAG and Visual-RAG-ME using GPT-4o as the LVLM reader and CLIP as the retriever. The original query results in low-quality retrieval augmentation, even slightly harming the performance on Visual-RAG in top-1 retrieval setting compared to no retrieval. By contrast, VisRet significantly improves QA accuracy in both top-1 and top-10 settings on both benchmarks, boosting accuracy to 0.538 on Visual-RAG and 0.700 on Visual-RAG-ME. Remarkably, on Visual-RAG-ME, VisRet outperforms top-10 retrieval in the original query setting with only top-1 retrieval, highlighting its high accuracy in retrieving the images containing the required features. Overall, the results confirm that VisRet not only improves retrieval accuracy but also leads to tangible gains in downstream VQA performance. In Section B.3, we further demonstrate that VisRet can bring similar performance gains to other models as the VQA reader. In Section B.4, we analyze the performance of directly using the generated images as the context and find that while T2I generation improves retrieval, it still cannot replace the retrieved natural images in most cases.
@@ 5 Conclusion
This work introduces VisRet, a framework that visualizes text queries to enable more accurate T2I retrieval. By operating entirely in the visual domain during retrieval, VisRet addresses key limitations of cross-modal embedding alignment. Our experiments confirm that visualized queries substantially improve both retrieval precision and downstream VQA accuracy across three benchmarking datasets and two retrievers. The simplicity and modularity of VisRet open up promising directions for future knowledge-intensive multi-modal systems.
@@
Although the goal of this paper is to introduce techniques to improve the text-to-image retrieval performance, the new approach could create new social risks. Specifically, in addition to the embedding model, our approach involves two neural networks: an LLM and a T2I generation model. It is possible for these large models to bring in new social bias in generating the visualized query and thus bias the retrieval results. For instance, when depicting certain scenes of social activity, T2I models could reinforce stereotypical social roles (Wan and Chang, 2024; Wan et al., 2024a). We urge practitioners to implement model debiasing and bias detection measure when deploying our proposed T2I retrieval method in real-world applications. Additionally, previous work has revealed problems with faithfulness and factuality in T2I models (Wan et al., 2024b; tse Huang et al., 2025; Wan and Chang, 2025). We also encourage future studies to explore solutions to similar trustworthy issue with multi-modal generative models.
@@ Appendix A Benchmark Data Details
In this section, we present the details of Visual-RAG data processing and Visual-RAG-ME annotation. Then, we provide the details of how we processed INQUIRE-Rerenk-Hard. Finally, we document the dataset statistics and the baseline T2I retrieval performance of CLIP (Radford et al., 2021) and E5-V (Jiang et al., 2024) in Table 3.
A.1 Visual-RAG and Visual-RAG-ME
Visual-RAG (Wu et al., 2025) releases 391 queries with associated image names from iNaturalist 2021 (Horn et al., 2021) and corresponding retrieval labels [^1] https://github.com/visual-rag/visual-rag. To prepare the data, we download the original iNaturalist 2021 dataset and re-collect the images from the train and test set. We were able to identify all annotated images in Visual-RAG except a single image due to a likely path error.
We annotate Visual-RAG-ME as an extension of Visual-RAG to visual question answering in the multi-entity setting. Concretely, Visual-RAG-ME reuses the visual features queried in Visual-RAG and constructs questions that compares those features between the organism covered in Visual-RAG with another similar entity, for which we manually annotate a new set of retrieval labels. The Visual-RAG-ME annotation pipeline consists of three steps: second entity identification; query composition and filtering; retrieval label annotation and balancing. We next describe the steps in detail.
Second Entity Identification
The goal of this step is to identify entities that are biologically close to the original entities in Visual-RAG so that plausible and challenging questions could be constructed. For this purpose, for each entity covered in Visual-RAG, we use BM25 (Robertson and Walker, 1994) to retrieve ten entities that have the closest full scientific names.
Query Composition and Filtering
In this step, we (the authors) manually traverse all the 391 questions in Visual-RAG and attempt to construct a corresponding multi-entity question. A question is constructed when we can identify images for the second entity that clearly depict the same feature as in the positive images for the original entity in Visual-RAG. The question we compose generally take a comparison style that asks whether the two organisms have the same feature or which organism feature a more extreme stylistic feature (e.g., lighter coloration, smoother surface etc.). We were able to construct 82 multi-entity questions after this step. Next, we perform a round of filtering to remove (1) the questions that both GPT-4o and GPT-4o-mini can answer correctly without image information and (2) questions can cover overly similar topics. After the filtering step, we ended up with 50 high quality multi-entity queries.
Retrieval Label Annotation and Balancing
Finally, for each question, we collect images of the second entity from iNaturalist and annotate their retrieval label. A positive label is assigned only if the image clearly displays the feature required to answer the question. For some of the questions, we find that a large number of positive images exist in the iNaturalist database. We therefore implement a filtering step where at most 50 positive images are kept for each entity. Table 8 shows two examples with their questions and ground-truth images.
In Table 3, we present the basic statistics of Visual-RAG-ME. While it has slightly more positive per entity compared to Visual-RAG, our new benchmark is still challenging, with both CLIP and E5-V achieving a low Recall@1 due to its lengthy and knowledge-intensive queries.
VR VR-ME IR-Hard
# Queries 391 50 59
||||Query|||| (word count) 18.5 25.1 6.0
# Images (per entity) 264 263 100
# Positives (per entity) 14.3 20.8 12.5
CLIP Recall@1 0.210 0.220 0.000
E5-V Recall@1 0.240 0.340 0.000
Table 3: Dataset statistics and baseline performance. VR denotes Visual-RAG and IR denotes INQUIRE-Rerank.
Method R@1 R@10 R@30 N@1 N@10 N@30
Baselines
Original Query 0.210 0.583 0.737 0.210 0.355 0.385
LLM Rephrase 0.238 0.586 0.737 0.238 0.360 0.395
VisRet
DALL-E 3 0.169 0.581 0.757 0.169 0.344 0.376
Stable Diffusion 3 0.166 0.517 0.691 0.166 0.319 0.349
Image-1 (low quality) 0.243 0.611 0.760 0.243 0.397 0.415
Image-1 (high quality) 0.251 0.645 0.793 0.251 0.431 0.438
Table 4: T2I Retrieval performance across different T2I generation models used for VisRet. We use GPT-4o to generate the T2I instruction and CLIP as the retriever. R = Recall. N = NDCG. The best results are boldfaced.
Retrieval Strategy LLM R@1 R@10 R@30 N@1 N@10 N@30
Original Query - 0.210 0.583 0.737 0.210 0.355 0.385
[…]
Llama 3.3 70B Instruct 0.240 0.575 0.742 0.240 0.377 0.399
GPT-4o 0.238 0.586 0.737 0.238 0.360 0.395
[…]
Llama 3.3 70B Instruct 0.256 0.627 0.790 0.256 0.413 0.437
GPT-4o 0.251 0.645 0.793 0.251 0.431 0.438
Table 5: Retrieval performance on Visual-RAG of with CLIP retriever, using LLMs as T2I instruction generator for VisRet. R = Recall. N = NDCG. The best results are boldfaced.
@@ A.2 INQUIRE-Rerank-Hard
To prepare INQUIRE-Rerank-Hard, we accessed the publicly released INQUIRE-Rerank (Vendrow et al., 2024) benchmark [^2] https://huggingface.co/datasets/evendrow/INQUIRE-Rerank. The original test set contained 160 queries, each paired with 100 images retrieved by CLIP. In a pilot study, we tested the retrieval performance of off-the-shelf CLIP and E5-V models. Results showed that CLIP can achieve 0.438 Recall@1 while E5-V achieved 0.506 Recall@1. After manually inspecting the data, we found that for a lot of instances, the negative images are not challenging enough and it is often very straightforward to identify the ground truth images. To highlight the challenging questions, we therefore filtered out the questions on which either CLIP and E5-V can achieve a perfect Recall@1. Overall, we observe that the remaining 59 questions require more nuanced image context understanding and a higher level of knowledge of the organism themselves, with more challenging confounder negative images.
Knowledge Visual-RAG Visual-RAG-ME
Model Knowledge Only 0 38.49 48.47 49.23 0 41.00 51.00 47.00
[…]
10 46.04 51.79 57.06 10 48.00 63.00 65.00
[…]
10 46.42 53.84 56.65 10 55.00 70.00 71.00
Knowledge Visual-RAG Visual-RAG-ME
Model Knowledge Only 0 38.49 48.47 49.23 0 41.00 51.00 47.00
[…]
10 46.42 53.84 56.65 10 55.00 70.00 71.00
Appendix B VisRet: Further Analyses
In this section, we provide more comprehensive analyses to investigate the effectiveness of VisRet from more perspectives, including the choices of T2I generation Model, T2I Instruction LLM, the downstream VQA LVLM reader. Finally, inspired by the generative retrieval literature, we conduct a pilot study of whether the generated images could be directly used as the knowledge context for downstream question answering.
B.1 T2I Generation Model Choice
How strong does the T2I generation model need to be for VisRet to work well? We compare the default T2I generation model (Image-1 with high quality setting) with three other models: DALL-E 3 (OpenAI, 2023a), Stable Diffusion 3 (Esser et al., 2024), and Image-1 with the low generation quality setting. Table 4 shows the results with GPT-4o as T2I instruction generation model and CLIP as the retriever model. Interestingly, compared to the cross-modality retrieval baselines, we find that DALL-E 3 and Stable Diffusion 3 do not provide significant performance improvements, while Image-1 low quality clearly and consistently improve the performance. The best performance is achieved by the newest and the most expensive Image-1 high quality setting. Together, these results suggest that a good T2I generation model with strong instruction following ability is necessary for VisRet. As further T2I generation methods improve, we anticipate that building more cost-efficient version of VisRet is a viable and promising further direction.
B.2 T2I Instruction LLM Choice
Does VisRet work well with other LLMs as the T2I instruction generator? In Table 5, we study two differently sized open-weight LLMs for rephrasing the query and generating the T2I instruction: Llama 3.1 8B Instruct and Llama 3.3 70B Instruct (Grattafiori et al., 2024). Overall, we observe promising results. For all the three LLMs, using them to generate T2I instructions for VisRet outperforms using the LLM themselves to rephrase the query. While more expensive LLMs achieve a high performance, the small 8B Llama model can already achieve decent performance at a similar level as GPT-4o.
B.3 Downstream VQA LVLM Choice
While we have shown the benefit of VisRet on VQA for GPT-4o, does the improvement hold across LVLMs with different capabilities? In Table 6, we repeat the VQA experiments with two additional LVLMs: GPT-4o-mini (version gpt-4o-mini-2024-07-18) and GPT-4.1 (version gpt-4.1-2025-04-14). Overall, we observe similar trends as those presented in Figure 2. Both direct T2I retrieval and VisRet outperform only relying on the model’s knowledge, with VisRet substantially outperforming the former. These results form the foundation for VisRet as a general plug-and-play method to enhance RAG pipelines that rely on accurate T2I retrieval.
B.4 Image Queries as Knowledge
As demonstrated by previous results, a T2I generation model with strong ability to follow instructions and generate realistic images is crucial to the success of VisRet. It is a natural question then, that it is still necessary to perform retrieval instead of directly using the generated images as the knowledge? In Table 7, we compare the performance of using a single image as the context with VisRet. Overall, we observe a mixed result. For Visual-RAG, GPT-4o-mini achieves slightly higher performance with the generated image than top-1 retrieval, but GPT-4o and GPT-4.1 exhibit the reverse pattern. For Visual-RAG-ME, both GPT-4o-mini and GPT-4.1 prefers the generated image over top-1 retrieval (and even top-10 retrieval). However, when provided with top-10 retrieved images, the models generally exhibit a higher VQA performance than using the generated image. Therefore, we conclude that retrieving natural images is still crucial for challenging VQA tasks like Visual-RAG and Visual-RAG-MR and cannot be fully replaced by pure T2I generation at this stage. It is an important future work to further combine image generation and image retrieval to improve the quality of the retrieved knowledge.
B.5 Further Qualitative Studies
In Table 8, we additionally provide some qualitative examples of how VisRet successfully improves T2I retrieval performance. As can be observed in top-ranked examples in text query-only retrieval, these retrieved images often fall short in correctly conveying visual semantics information such as angle, body part depicted, form of the subject depicted, visual distance of the subject, and so on. In contrast, VisRet is able to represent the inferred nuanced visual semantics in the image generation step first, then utilize this within-modality semantic-rich image query to obtain precise visual knowledge related to the VQA task.
Appendix C Implementation Details

### v2

@@ 5.1 Experimental Setup
We evaluate on four challenging benchmarks: (1) Visual-RAG (Wu et al., 2025b), a T2I retrieval and VQA benchmark featuring visual knowledgeintensive questions on features of natural species that are not commonly documented in text corpora. (2) Visual-RAG-ME, a new benchmark we introduce featuring queries that compare the same visual feature across multiple entities. (3) INQUIRE-Rerank (Vendrow et al., 2024), a T2I retrieval benchmark requiring accurate knowledge of species appearance and behavior. (4) Microsoft COCO 2017 (Chen et al., 2015), a widely used image captioning dataset, which we repurpose for caption-to-image retrieval evaluation. For (3) and (4), we perform additional filtering to retain the most difficult queries and call the resulting datasets INQUIRE-Rerank-Hard and COCO-Hard. Table˜1 shows the basic statistics of each dataset. We present additional details in Appendix Appendix˜A.
Metrics
@@
For all the three benchmarks, we evaluate T2I retrieval with Recall@$k$ and nDCG@$k$ with $k$ = 1, 10, 30111As nDCG@1 is undefined, we use Recall@1 as its value.. For Visual-RAG and Visual-RAG-ME, we additionally use an LLM judge to evaluate the end-to-end VQA accuracy, using a prompt derived from Visual-RAG (Wu et al., 2025b).
Baselines
Using the benchmarks, we evaluate three families of T2I retrieval methods:
  1. 1.
    Text-to-Image Matching: Retrieving with the original textual query against the image index using off-the-shelf retrievers (CLIP, E5-V). We also compare with a baseline of applying query rewriting via an LLM to highlight the desired features and then performing standard T2I retrieval over the same image embedding index.
  2. 2.
    Text-to-Text Matching: Converting the images into textual surrogates for text-to-text retrieval. We compare with (1) captioning the entire corpus with BLIP, (2) top-k reranking with VISA, and (3) an ablated version of VISA, where we rerank top-k images using LLM-generated captions. We use Qwen2.5-70B-VL (Bai et al., 2025) as the model for (2) and (3) and rerank the top-30 retrieval.
  3. 3.
    Image-to-Image Matching: We evaluate VisRet and an image ablation of VisRet that only uses a single image instead of multiple images (-multi-image).
Implementation Details
For all the retrieval experiments, we use CLIP and E5-V as the retrievers, GPT-4o (OpenAI, 2023b) as T2I instruction generation LLM, and gpt-image-1 (OpenAI, 2025) as the T2I Model to generate $m=3$ images. For the downstream question answering (Visual-RAG and Visual-RAG-ME), we use GPT-4o as the reader LLM. We present the full implementation details of VisRet and the baselines in Appendix˜B.
5.2 VisRet Improves T2I Retrieval
[…]
Compared with VisRet, the text-to-text retrieval baselines exhibit inconsistent performance across the benchmarks. Captioning the entire corpus with BLIP only works on COCO, which has the closest relationship to its training data, while it does not improve performance over cross-modal retrieval on the other benchmarks with more knowledge intensive queries. We hypothesize that this is caused by the high information density in images, which incurs an inevitable loss of potentially useful information during captioning. On the other hand, with access to cross-modal retrieval results and by directly operating on them, Captioning Reranking and VISA outperform query-agnostic corpus captioning, yet cannot outperform direct cross-modal retrieval by a large margin. Furthermore, these methods can only process a limited number of top-k retrieved items as their cost grows linearly with k. As a result, their performance is limited by the first round of retrieval, leading to near-zero performance on INQUIRE-Rerank-Hard and COCO-Hard.
Finally, comparing VisRet with its single-image variant, we find that using only one generated image as the query slightly harms the performance but still outperforms most of the other baselines. The finding holds for both the lightweight CLIP embedding or the stronger E5-V embeddings, indicating the flexibility of VisRet. Table˜3 presents several examples to demonstrate how the visualization step successfully captures subtle visual features implied by the original text query. As shown in the examples, for a number of ground truth images, the main entity is either in a particular posture that reveals the required information (Visual-RAG and Visual-RAG-ME) or in a complex relationships between other entities (INQUIRE and COCO), posing a challenge to cross-modal similarity matching to recognize and satisfy these constraints. On the other hand, VisRet infuses these constraints naturally into the image query and significantly reduces the workload of the downstream embedding model.
@@ 5.2 VisRet Improves T2I Retrieval
5.3 VisRet improves Downstream QA
Retrieval Strategy LLM R@1 R@10 R@30 N@1 N@10 N@30
Original Query - 0.210 0.583 0.737 0.210 0.355 0.385
LLM Rewriting Llama 3.1 8B Instruct 0.238 0.563 0.737 0.238 0.365 0.385
Llama 3.3 70B Instruct 0.240 0.575 0.742 0.240 0.377 0.399
GPT-4o 0.238 0.586 0.737 0.238 0.360 0.395
Visualize-then-Retrieve Llama 3.1 8B Instruct 0.243 0.606 0.780 0.243 0.405 0.428
Llama 3.3 70B Instruct 0.256 0.627 0.790 0.256 0.413 0.437
GPT-4o 0.251 0.645 0.793 0.251 0.431 0.438
Table 4: Retrieval performance on Visual-RAG of with CLIP retriever, using different LLMs as T2I instruction generator for VisRet. R = Recall. N = nDCG. The best results are boldfaced.
@@ 5.3 VisRet improves Downstream QA
To assess the utility of VisRet in real-world applications, we evaluate its downstream VQA accuracy in a RAG pipeline. We compare three settings: (1) using only the model’s internal knowledge, (2) RAG with original text query-based retrieval and (3) RAG with VisRet. Figure˜2 shows the QA accuracy on Visual-RAG and Visual-RAG-ME using GPT-4o as the LVLM reader and CLIP as the retriever. The original query results in low-quality retrieval augmentation, even slightly harming performance on Visual-RAG in top-1 retrieval setting compared to no retrieval. By contrast, VisRet significantly improves QA accuracy in both top-1 and top-10 settings on both benchmarks, boosting accuracy to 0.538 on Visual-RAG and 0.700 on Visual-RAG-ME. Remarkably, on Visual-RAG-ME, VisRet outperforms top-10 retrieval in the original query setting with only top-1 retrieval, highlighting its high accuracy in retrieving the images that contain the required features. Overall, the results confirm that VisRet not only improves retrieval accuracy but also leads to tangible gains in downstream VQA performance. In Section˜6.3, we further demonstrate that VisRet can bring similar performance gains to other models as the VQA reader. In Section˜6.4, we analyze the performance of directly using the generated images as the context and find that while T2I generation improves retrieval, it still cannot replace the retrieved natural images in most cases.
6 Analyses
In this section, we provide more comprehensive analyses to investigate the effectiveness of VisRet from more perspectives, including the choices of T2I instruction LLM, T2I generation model, and the downstream VQA reader. Finally, inspired by the generative retrieval literature, we conduct a pilot study of whether the generated images could be directly used as the knowledge context for downstream question answering.
6.1 Query Generation Model Choice
Does VisRet work well with other LLMs as the T2I instruction generator? In Table˜4, we study two differently sized open-weight LLMs for rephrasing the query and generating the T2I instruction: Llama 3.1 8B Instruct and Llama 3.3 70B Instruct (Grattafiori et al., 2024). Overall, we observe promising results: For all three LLMs, using the LLMs themselves to generate T2I instructions for VisRet outperforms using the LLM themselves to rephrase the query. While more expensive LLMs achieve higher performance, the small 8B Llama model can already achieve decent performance at a similar level as GPT-4o.
Method N@1 N@10 N@30 Avg
Retriever = CLIP
Original Query 0.220 0.423 0.435 0.359
VisRet Variants:
Stable Diffusion 3.5 0.270 0.467 0.484 0.407
FLUX.1-dev 0.320 0.501 0.494 0.438
Qwen-Image 0.330 0.501 0.518 0.450
DALL-E 3 0.346 0.554 0.553 0.484
Gemini 0.410 0.579 0.583 0.524
Image-1 (low quality) 0.420 0.585 0.589 0.531
Image-1 (high quality) 0.460 0.632 0.605 0.566
Retriever = E5-V
Original Query 0.340 0.465 0.486 0.430
VisRet Variants:
Stable Diffusion 3.5 0.410 0.550 0.547 0.502
FLUX.1-dev 0.370 0.535 0.532 0.479
Qwen-Image 0.410 0.542 0.548 0.500
DALL-E 3 0.450 0.599 0.587 0.545
Gemini 0.615 0.585 0.666 0.589
Image-1 (low quality) 0.520 0.629 0.612 0.587
Image-1 (high quality) 0.560 0.643 0.622 0.608
Table 5: nDCG@k retrieval results on Visual-RAG-ME for different T2I generation models used in VisRet. We report N@1, N@10, N@30, and their average (Avg) for two retrievers (CLIP and E5-V). Best numbers within each retriever are boldfaced. Gemini stands for the Gemini-2.5-Flash-Image-Preview model.
Knowledge Visual-RAG Visual-RAG-ME
Model Knowledge Only 0 0.385 0.485 0.492 0 0.410 0.510 0.470
Direct T2I Retrieval 1 0.409 0.474 0.515 2 0.490 0.590 0.610
10 0.460 0.518 0.571 10 0.480 0.630 0.650
Visualize-then-Retrieve 1 0.418 0.492 0.572 2 0.530 0.640 0.620
10 0.464 0.538 0.567 10 0.550 0.700 0.710
Knowledge Visual-RAG Visual-RAG-ME
Model Knowledge Only 0 0.385 0.485 0.492 0 0.410 0.510 0.470
Generated Image (Image-1) 1 0.431 0.425 0.444 2 0.590 0.580 0.800
Visualize-then-Retrieve 1 0.418 0.492 0.572 2 0.530 0.640 0.620
10 0.464 0.538 0.567 10 0.550 0.700 0.710
6.2 T2I Generation Model Choice
How strong does the T2I generation model need to be for VisRet to work well? We compare the default T2I generation model (Image-1 with high quality setting) with several other commonly used strong T2I generation models:
  • Stable Diffusion 3 (Esser et al., 2024).
  • FLUX-1.dev (Black Forest Labs, 2024).
  • Qwen-Image (Wu et al., 2025a).
  • DALL-E 3 API (OpenAI, 2023a).
  • Gemini-2.5-Flash-Image-Preview API (Fortin et al., 2025).
  • Image-1 API with the low quality setting.
Table˜5 shows the results with GPT-4o as the T2I instruction generation model and CLIP as the retriever model. We observe that while all the models can improve over cross-modality retrieval to some extent in the VisRet pipeline, their effectiveness increases as the image quality retrieves. The best performance is achieved by the latest proprietary models Image-1 and Gemini-2.5-Flash-Image-Preview. Together, these results suggest that a good T2I generation model with strong instruction following ability is necessary for VisRet. As further T2I generation methods improve, we anticipate that building more cost-efficient versions of VisRet is viable and promising.
6.3 Downstream VQA Model Choice
While we have shown the benefit of VisRet on VQA for GPT-4o, does the improvement hold across LVLMs with different capabilities? In Table˜6, we repeat the VQA experiments with two additional LVLMs: GPT-4o-mini (version gpt-4o-mini-2024-07-18) and GPT-4.1 (version gpt-4.1-2025-04-14). Overall, we observe similar trends as those presented in Figure˜2. Both direct T2I retrieval and VisRet outperform only relying on the model’s knowledge, with VisRet substantially outperforming the former. These results form the foundation for VisRet as a general plug-and-play method to enhance RAG pipelines that rely on accurate T2I retrieval.
6.4 Image Queries as Knowledge
As demonstrated by previous results, a T2I generation model with strong ability to follow instructions and generate realistic images is crucial to the success of VisRet. A natural question is whether it is still necessary to perform retrieval instead of directly using the generated images as the knowledge? In Table˜7, we compare the performance of using a single image as the context with VisRet. Overall, we observe mixed results. For Visual-RAG, GPT-4o-mini achieves slightly higher performance with the generated image than top-1 retrieval, but GPT-4o and GPT-4.1 exhibit the reverse pattern. For Visual-RAG-ME, both GPT-4o-mini and GPT-4.1 prefer the generated image over top-1 retrieval (and even top-10 retrieval). However, when provided with top-10 retrieved images, the models generally exhibit a higher VQA performance than when using the generated image. Therefore, we conclude that retrieving natural images is still crucial for challenging VQA tasks like Visual-RAG and Visual-RAG-ME and cannot be fully replaced by pure T2I generation at this stage. An important further direction is to combine image generation and image retrieval to improve the quality of the retrieved knowledge.
@@ 7 Conclusion
This work introduces VisRet, a novel framework that visualizes text queries to reframe cross-modality T2I retrieval as within-modality retrieval. By operating entirely in the visual domain, VisRet addresses key limitations of cross-modal embedding alignment. Our experiments confirm that visualized queries substantially improve both retrieval precision and downstream VQA accuracy across four benchmark datasets and two retrievers. The simplicity and modularity of VisRet open up promising directions for future knowledge-intensive multimodal systems.
@@
Although the goal of this paper is to introduce techniques to improve the text-to-image retrieval performance, the new approach could create new social risks. Specifically, in addition to the neural embedding model, our approach involves two neural models: an LLM and a T2I generation model. It is possible for these large models to bring in new social bias in generating the visualize query and thus bias the retrieval results. For instance, when depicting certain scenes of social activity, the models could reinforce stereotypical social roles. We urge practitioners to implement model debiasing and bias detection measure when deploying our proposed T2I retrieval method in real-world applications.
@@
AI Assistant Use
AI assistants, specifically ChatGPT, are used only for revising the paper draft, fixing grammar mistakes, and improving the outlook of the figures.
@@ Appendix A Benchmark Data Details
In this section, we present the details of data processing for Visual-RAG, INQUIRE-Rerenk-Hard, and COCO-Hard.
A.1 Visual-RAG
Visual-RAG (Wu et al., 2025b) releases 391 queries with associated image names from iNaturalist 2021 (Horn et al., 2021) and corresponding retrieval labels [^2] https://github.com/visual-rag/visual-rag. To prepare the data, we download the original iNaturalist 2021 dataset and re-collect the images from the train and test set. We were able to identify all annotated images in Visual-RAG except a single image due to a likely path error.
@@ A.2 INQUIRE-Rerank-Hard
To prepare INQUIRE-Rerank-Hard, we accessed the publicly released INQUIRE-Rerank (Vendrow et al., 2024) benchmark [^3] https://huggingface.co/datasets/evendrow/INQUIRE-Rerank. The original test set contained 160 queries, each paired with 100 images retrieved by CLIP. In a pilot study, we tested the retrieval performance of off-the-shelf CLIP and E5-V models. Results showed that CLIP can achieve 0.438 Recall@1 while E5-V achieved 0.506 Recall@1. After manually inspecting the data, we found that for a lot of instances, the negative images are not challenging enough and it is often very straightforward to identify the ground truth images. To highlight the challenging questions, we therefore filtered out the questions on which either CLIP or E5-V can achieve a perfect Recall@1. Overall, we observe that the remaining 59 questions require more nuanced image context understanding and a higher level of knowledge of the organisms themselves, with more challenging confounder negative images.
A.3 COCO-Hard
We derive COCO-Hard from the publicly released validation set of Microsoft COCO 2017 (Chen et al., 2015). The original validation set contained 5000 queries, and we directly use the ground truth images themselves to construct the image corpus. In a pilot study, we found that CLIP is able to achieve 0.424 Recall@1 and 0.756 Recall@10, while E5-V has 0.594 Recall@1 and 0.892 Recall@10. A manual analysis shows that the queries of most of the examples in the original COCO dataset are short phrases centered on a single entity. To increase the task’s difficulty, we rank the examples in reverse order by nDCG@30 and retained the hardest 500 instances. We find these hard queries often describe complex spatial relationships among multiple entities.
Appendix B Implementation Details